The two upper front teeth of a person exert a total force of 30 N on a piece of steak. If
the biting edge of each tooth measure 10 mm x 1.0 mm, what is the pressure oe steak?
1
Expert's answer
2021-07-19T09:52:16-0400

The pressure

"P=\frac{F}{A}=\frac{30\:\rm N}{10*10*10^{-6}\:\rm m^2}=300\: 000\:\rm Pa"